OpenStudy (anonymous):

Find the rate of change of the area A of a circle with respect to its circumference C.

OpenStudy (anonymous):

A = pi r^2 C = 2pi*r --> r = C / (2*pi) Substitute: A = C^2 /4 pi dA/dC = C / (2 pi)